OpenStudy (blackops2luvr):

3/-4 = y/-16 what value of y makes the equation true?

OpenStudy (blackops2luvr):

\(\Huge {\color{blue} {help} }\) \(\Huge {\color{orange} {anyone?} }\)

OpenStudy (austinl):

\(\displaystyle \frac{3}{-4}=\frac{y}{-16}\) Multiply both sides by -16.

OpenStudy (blackops2luvr):

..

OpenStudy (austinl):

\(\displaystyle\frac{3}{-4}\times\frac{-16}{1}=~?\)

OpenStudy (blackops2luvr):

12

OpenStudy (blackops2luvr):

\(\Huge {\color{blue} {help} }\) \(\Huge {\color{orange} {anyone?} }\)

OpenStudy (phi):

can you simplify 12/-16 ?

OpenStudy (phi):

what do you get if you divide top and bottom by 4 ?